# Involutive (simple) latin solutions of the Yang-Baxter equation and related (left) quasigroups

### v1

@@ 3.2 Indecomposable cycle sets of prime-power size
The converse follows by [21, Lemma 4.5]. ∎
@@
The previous result reduces the description of all indecomposable and irretractable cycle sets of prime-power size $p^{n}$ and with normal Sylow $p$-subgroup permutation group to the ones having a $p$-group permutation group. This is consistent with the first part of [21, Section $4$]. However, we highlight that 22 is not useful to describe all the irretractable and indecomposable cycle set of size $p^{n}$. Indeed, Example $5$ of [40] is an irretractable and indecomposable cycle set of size $8$ and with permutation group isomorphic to $\mathbb{S}_{4}$.
@@
It follows by 22 and [21, Lemma 4.5]. ∎
As a corollary, we recover [24, Theorem 2.13], that classify indecomposable cycle sets of prime size.
@@
Suppose that $X$ is irretractable. Since by [24, Theorem $2.15$] $\mathcal{G}(X)$ is solvable, by [32, Exercise $7.2.12$] we have that it is isomorphic to a subgroup of the form $\mathbb{Z}/p\mathbb{Z}\rtimes N$, with $N$ a cyclic subgroup having order dividing $p-1$. Therefore, by 22 $X$ is a deformation of an indecomposable cycle set $\bar{X}$ with permutation group isomorphic to $\mathbb{Z}/p\mathbb{Z}$. Since $X$ is irretractable, it follows that $\bar{X}$ is irretractable, but this contradicts [13, Proposition 1]. Hence, we have that $X$ is retractable. Therefore by [11, Lemma 1] $X$ is a cycle set given by $x\cdotp y:=\alpha(y)$ for all $x,y\in X$, where $\alpha$ is a $p$-cycle. By a standard calculation, one can show that the isomorphism class of this cycle set does not depend on the choice of the $p$-cycle. ∎
@@ 3.3 Blocks of imprimitivity
At first, recall that a permutation group $G$ of a finite set $X$ is said to be quasiprimitive if every normal subgroup acts transitively on $X$ and it is called innately transitive if there exist a minimal normal subgroup that acts transitively on $X$. Clearly, the implications “primitive $\Rightarrow$ quasiprimitive” and “quasiprimitive $\Rightarrow$ innately transitive” hold. For solvable groups, these notions are equivalent (for more details, see [5, Section 1] and [30]).
@@
By [24, Theorem 2.15], $\mathcal{G}(X)$ is a finite solvable group, hence by Lemma 25 we have that $\mathcal{G}(X)$ acts primitively on $X$. Again by solvability of $\mathcal{G}(X)$ and by [32, Theorem 7.2.6] we have that $|X|=p^{n}$ for some natural number $n$ and $\mathcal{G}(X)$ is isomorphic to the semidirect product $N\rtimes G$ where $N$ is an elementary abelian $p$-group that acts transitively on $X$ and $G$ is the stabilizer of an element $x\in X$. Moreover, $N$ is a minimal normal subgroup of $\mathcal{G}(X)$ and is self-centralizing. By primitivity of $\mathcal{G}(X)$, either $X$ is irretractable or $n$ is equal to $1$. Suppose that $X$ is irretractable. Since $\mathcal{G}(X)$ is solvable and finite, we can suppose that $G$ contains $B_{p_{2}}+...+B_{p_{r}}$ and by 19, we have that $X$ can be identified with a transitive cycle base, therefore $\lambda_{a}(x)=x$ for all $a\in G$ and, if $x_{1}\in B_{p_{1}}$,…,$x_{r}\in B_{p_{r}}$ are such that $x=x_{1}+...+x_{r}$, it follows that $\lambda_{a}(x_{i})=x_{i}$ for all $i\in\{1,...,r\}$. Moreover, since $N$ is a normal $p$-group, by Lemma 16 it follows that $\lambda_{n}(x_{i})=x_{i}$ for all $i\in\{2,...r\}$. By these facts, we obtain that $B_{p_{2}}+...+B_{p_{r}}$ is a trivial left brace with cyclic additive group, and that $B_{p}$ acts trivially on $B_{p_{2}}+...+B_{p_{r}}$ by the map $\lambda$, therefore by Lemma 16 $B_{p}$ is normal in $\mathcal{G}(X)$. By minimality and self-centrality of $N$, it follows that $N=Z(B_{p})=B_{p}$. By 22, $X$ is a deformation of an irretractable cycle set $\bar{X}$ with abelian permutation group isomorphic to $N$, but this contradicts [13, Proposition 1]. Therefore, $X$ has prime size $p$. The rest of the statement follows by 24. ∎
@@ 4.1 Right linear cycle sets
([22, Proposition $3.9$]) Let $X=\mathrm{Aff}(G,\phi,\psi,c)$ be a latin left quasigroup affine over an abelian group $(G,+)$. Then
@@
    If $X$ is a cycle set, then $X$ is right linear over $G$ if and only if $X_{\alpha}$ is right linear over $G$.
@@
4) By 1), 2) and [8, Theorem $5.3$], we have that if $X$ is a cycle set right linear over a group $G$, then so is $X_{\alpha}$. The converse follows as in 3). ∎
@@
We indicate by $\sim$ the equivalence relation induced by $\mathcal{B}$. At first, we show that $B$ is a subgroup of $G$. Now, let $y_{1},y_{2}\in B$. Then, since $\phi$ is a bijection, we have that $s\circ\psi(y_{1})\sim s\circ\psi(y_{2})$ for all $s\in G$. If we set $s:=\psi(y_{1})^{-}$, since $\psi\in Aut(G,\circ)$ we obtain $\psi(y_{1}^{-}\circ y_{2})\sim 0$, therefore $\psi(y_{1}^{-}\circ y_{2})\in B$ and, hence $y_{1}^{-}\circ y_{2}\in B$. Therefore, it follows that $B$ is a subgroup of $G$. By a standard calculation, we obtain that $\mathcal{B}=\mathcal{B}_{B}$, hence the statement follows. ∎
The following result improves [8, Lemma 4.8] in the finite case.
@@ 4.2 Right linear representation
A left quasigroup $X$ has semiregular displacement group if and only if $\sim_{\mathrm{Dis}(X)}=1_{X}$ (see Eq. 3). Left quasigroups with semiregular displacement groups have been studied in [7], and in particular we have the following corollary of [7, Lemma 3.2].
@@
The connection between right linear left quasigrups and cycle sets has been studied in [8], in particular, we have the following reformulation of [8, Theorem $5.3$], since finite irretractable cycle sets with semiregular displacement group are latin.
@@
If $X$ is an irretractable cycle set, by a standard calculation we have that if either 1)1)1 ) or 2)2)2 ) hold, then $X$ is latin. Hence, the statement follows by [8, Theorem 5.3]. ∎
@@
If $(X,\cdotp)$ is an irretractable cycle set with regular displacement group, by 33 we can represent it as a cycle set right linear over a suitable group $(G,\circ)$ and suitable maps $\phi$ and $\psi$ as in 5. The theory contained in [8, Section 5] gives a method to find the operation $\circ$ and the maps $\phi$ and $\psi$, even if these maps are not uniquely determined. Indeed, if we choose two elements $e,f\in X$, then $X$ has a group structure given by $x\circ_{e,f}y:=\delta_{e}^{-1}(x)\cdotp\sigma_{f}^{-1}(y)$ for all $x,y\in X$ , where $f\cdotp e$ is the neutral element. In this way, $\phi_{e,f}(x):=\delta_{e}(x)\circ_{e,f}\sigma_{f}(e\cdotp f)$ and $\psi_{e,f}(x):=\sigma_{f}(e\cdotp f)^{-1}\circ_{e,f}\sigma_{f}(x)$ for all $x\in X$. The displacement group $\mathrm{Dis}(X)$ is isomorphic to $(X,\circ_{e,f})$, and an explicit isomorphism is given by $\rho_{e,f}:(X,\circ_{e,f})\rightarrow\mathrm{Dis}(X)$, $x\mapsto\sigma_{\delta_{e}^{-1}(x)}\sigma_{f}^{-1}$ for all $x\in X$. In this way, the map $\psi_{e,f}$ induces an automorphism of $\mathrm{Dis}(X)$ given by $\overline{\psi_{e,f}}(\sigma_{\delta_{e}^{-1}(x)}\sigma_{f}^{-1}):=\sigma_{\delta_{e}^{-1}(\psi_{e,f}(x))}\sigma_{f}^{-1}$ and the map $\phi_{e,f}$ induces a bijection given by $\overline{\phi_{e,f}}(\sigma_{\delta_{e}^{-1}(x)}\sigma_{f}^{-1}):=\sigma_{\delta_{e}^{-1}(\phi_{e,f}(x))}\sigma_{f}^{-1}$, for all $x\in X$. Therefore, under the correspondence $\rho_{e,f}$, the underlying group $(X,\circ_{e,f})$ can be identified with the displacement group and without loss of generality we can say that $X$ is right linear over $\mathrm{Dis}(X)$.
@@
By 33, $X$ can be identified with a cycle set right linear over a group $(G,\circ)$ and given by $x\cdotp y:=\phi(x)\circ\psi(y)$ for all $x,y\in G$, for a suitable maps $\phi$ and $\psi$. By 31 the congruences of $X$ are in one-to-one correspondence with normal subgroups $H$ of $G$ that are invariant under $\phi$ and $\psi$. Since $\mathrm{Dis}(X)$ acts on $X$ by left translations (respect to the operation $\circ$), the equivalence classes of the congruence induced by a subgroup $H$ of $G$ correspond to the orbits respect to the action of the subgroup $T_{H}$ of $\mathrm{Dis}(X)$ given by the translations by the elements of $H$. Therefore $T_{H}$ is an ideal by [34, Theorem 1]. Clearly, every ideal of $\mathcal{G}(X)$ contained in $\mathrm{Dis}(X)$ induces a congruence given by its orbits. Since $\mathrm{Dis}(X)$ acts by left translation with respect to the operation $\circ$, the statement follows. ∎
@@ 5.1 Simple left quasigroups
As it is well-known that if $X$ is an indecomposable cycle set then the ideal $\mathcal{G}(X)^{2}$ of $\mathcal{G}(X)$ coincides with $\mathrm{Dis}(X)$ (see [10, Lemma 11]), we can use our results to recover a characterization of simple cycle sets. Since the unique simple decomposable cycle set has size $2$, we decide to add the condition $|X|>2$ and isolate this easy case.
@@
If $X$ is simple and is not the trivial cycle set of prime size, by [16, Proposition 4.1] it is irretractable, hence this implication follows by 20, 38 and the equality $\mathcal{G}(X)^{2}=\mathrm{Dis}(X)$. Conversely, if $X$ is the trivial indecomposable of prime size, then it is simple by [11, Lemma 1]. Now, suppose that $X$ is irretractable and $\mathcal{G}(X)^{2}$ is a minimal ideal of $\mathcal{G}(X)$ that acts transitively on $X$. Then $X$ is indecomposable and hence we have $\mathcal{G}(X)^{2}=\mathrm{Dis}(X)$. Therefore, the thesis follows by 20 and 38. ∎
@@ 5.2 Simple cycle sets with nilpotent displacement group
If $H$ is a non-trivial normal Hall subgroup of $\mathrm{Dis}(X)$, then by [3, Proposition 4.2.1] it is an ideal of $\mathrm{Dis}(X)$. Moreover, $H$ is a characteristic subgroup of $(\mathrm{Dis}(X),+)$ and $(\mathrm{Dis}(X),\circ)$, therefore it is an ideal of $\mathcal{G}(X)$. ∎
@@
1) $\Rightarrow$ 3) Suppose that $X$ is a simple cycle set with cyclic displacement group and such that $|X|$ is not a prime number. Then, $X$ must be irretractable and hence $\mathrm{Dis}(X)$ acts transitively on $X$, therefore by 41 we have $|X|=|\mathrm{Dis}(X)|=p^{k}$ for some prime number $p$. Moreover, by an inspection of the small cycle sets, $|X|\neq 4$. Now, let $m$ be a divisor of $|X|$ such that $|X|/m=p$ for a prime number $p$. By the results containted in [35] and [2, Proposition 5.4], the set $H:=p\mathrm{Dis}(X)$ is an ideal of $\mathrm{Dis}(X)$ of size $m$ and is a characteristic subgroup of $(\mathrm{Dis}(X),+)$ and $(\mathrm{Dis}(X),\circ)$, hence it is an ideal of $\mathcal{G}(X)$. Therefore, $p\mathrm{Dis}(X)$ induces a non-trivial congruence on $X$, a contradiction. Then, $|X|$ must be a prime number, and by [16, Proposition 4.1] $X$ is indecomposable or $|X|=2$. In both cases, by [24, Theorem 2.13] we have $\mathrm{Dis}(X)=\{id_{X}\}$
3) $\Rightarrow$ 2) If $X$ is a simple cycle set of prime size, we have that $\mathrm{Dis}(X)$ is the trivial group by [24, Theorem 2.13]

### v2

@@ 3.2 Indecomposable cycle sets of prime-power size
The converse follows by [22, Lemma 4.5]. ∎
@@
The previous result reduces the description of all indecomposable and irretractable cycle sets of prime-power size $p^{n}$ and with normal Sylow $p$-subgroup permutation group to the ones having a $p$-group permutation group. This is consistent with the first part of [22, Section $4$]. However, we highlight that 22 is not useful to describe all the irretractable and indecomposable cycle set of size $p^{n}$. Indeed, Example $5$ of [41] is an irretractable and indecomposable cycle set of size $8$ and with permutation group isomorphic to $\mathbb{S}_{4}$.
@@
It follows by 22 and [22, Lemma 4.5]. ∎
As a corollary, we recover [25, Theorem 2.13], that classify indecomposable cycle sets of prime size.
@@
Suppose that $X$ is irretractable. Since by [25, Theorem $2.15$] $\mathcal{G}(X)$ is solvable, by [33, Exercise $7.2.12$] we have that it is isomorphic to a subgroup of the form $\mathbb{Z}/p\mathbb{Z}\rtimes N$, with $N$ a cyclic subgroup having order dividing $p-1$. Therefore, by 22 $X$ is a deformation of an indecomposable cycle set $\bar{X}$ with permutation group isomorphic to $\mathbb{Z}/p\mathbb{Z}$. Since $X$ is irretractable, it follows that $\bar{X}$ is irretractable, but this contradicts [14, Proposition 1]. Hence, we have that $X$ is retractable. Therefore by [12, Lemma 1] $X$ is a cycle set given by $x\cdotp y:=\alpha(y)$ for all $x,y\in X$, where $\alpha$ is a $p$-cycle. By a standard calculation, one can show that the isomorphism class of this cycle set does not depend on the choice of the $p$-cycle. ∎
@@ 3.3 Blocks of imprimitivity
At first, recall that a permutation group $G$ of a finite set $X$ is said to be quasiprimitive if every normal subgroup acts transitively on $X$ and it is called innately transitive if there exist a minimal normal subgroup that acts transitively on $X$. Clearly, the implications “primitive $\Rightarrow$ quasiprimitive” and “quasiprimitive $\Rightarrow$ innately transitive” hold. For solvable groups, these notions are equivalent (for more details, see [6, Section 1] and [31]).
@@
By [25, Theorem 2.15], $\mathcal{G}(X)$ is a finite solvable group, hence by Lemma 25 we have that $\mathcal{G}(X)$ acts primitively on $X$. Again by solvability of $\mathcal{G}(X)$ and by [33, Theorem 7.2.6] we have that $|X|=p^{n}$ for some natural number $n$ and $\mathcal{G}(X)$ is isomorphic to the semidirect product $N\rtimes G$ where $N$ is an elementary abelian $p$-group that acts transitively on $X$ and $G$ is the stabilizer of an element $x\in X$. Moreover, $N$ is a minimal normal subgroup of $\mathcal{G}(X)$ and is self-centralizing. By primitivity of $\mathcal{G}(X)$, either $X$ is irretractable or $n$ is equal to $1$. Suppose that $X$ is irretractable. Since $\mathcal{G}(X)$ is solvable and finite, we can suppose that $G$ contains $B_{p_{2}}+...+B_{p_{r}}$ and by 19, we have that $X$ can be identified with a transitive cycle base, therefore $\lambda_{a}(x)=x$ for all $a\in G$ and, if $x_{1}\in B_{p_{1}}$,…,$x_{r}\in B_{p_{r}}$ are such that $x=x_{1}+...+x_{r}$, it follows that $\lambda_{a}(x_{i})=x_{i}$ for all $i\in\{1,...,r\}$. Moreover, since $N$ is a normal $p$-group, by Lemma 16 it follows that $\lambda_{n}(x_{i})=x_{i}$ for all $i\in\{2,...r\}$. By these facts, we obtain that $B_{p_{2}}+...+B_{p_{r}}$ is a trivial left brace with cyclic additive group, and that $B_{p}$ acts trivially on $B_{p_{2}}+...+B_{p_{r}}$ by the map $\lambda$, therefore by Lemma 16 $B_{p}$ is normal in $\mathcal{G}(X)$. By minimality and self-centrality of $N$, it follows that $N=Z(B_{p})=B_{p}$. By 22, $X$ is a deformation of an irretractable cycle set $\bar{X}$ with abelian permutation group isomorphic to $N$, but this contradicts [14, Proposition 1]. Therefore, $X$ has prime size $p$. The rest of the statement follows by 24. ∎
@@ 4.1 Right linear cycle sets
([23, Proposition $3.9$]) Let $X=\mathrm{Aff}(G,\phi,\psi,c)$ be a latin left quasigroup affine over an abelian group $(G,+)$. Then
@@
    If $X$ is a latin cycle set, then $X$ is right linear over $G$ if and only if $X_{\alpha}$ is right linear over $G$.
@@
4) By 1), 2) and [9, Theorem $5.3$], we have that if $X$ is a latin cycle set right linear over a group $G$, then so is $X_{\alpha}$. The converse follows as in 3). ∎
@@
We indicate by $\sim$ the equivalence relation induced by $\mathcal{B}$. At first, we show that $B$ is a subgroup of $G$. Now, let $y_{1},y_{2}\in B$. Then, since $\phi$ is a bijection, we have that $s\circ\psi(y_{1})\sim s\circ\psi(y_{2})$ for all $s\in G$. If we set $s:=\psi(y_{1})^{-}$, since $\psi\in Aut(G,\circ)$ we obtain $\psi(y_{1}^{-}\circ y_{2})\sim 0$, therefore $\psi(y_{1}^{-}\circ y_{2})\in B$, and hence $y_{1}^{-}\circ y_{2}\in B$. Therefore, it follows that $B$ is a subgroup of $G$. By a standard calculation, we obtain that $\mathcal{B}=\mathcal{B}_{B}$, hence the statement follows. ∎
The following result improves [9, Lemma 4.8] in the finite case.
@@ 4.2 Right linear representation
A left quasigroup $X$ has semiregular displacement group if and only if $\sim_{\mathrm{Dis}(X)}=1_{X}$ (see Eq. 3). Left quasigroups with semiregular displacement groups have been studied in [8], and in particular we have the following corollary of [8, Lemma 3.2].
@@
The connection between right linear left quasigrups and cycle sets has been studied in [9], in particular, we have the following reformulation of [9, Theorem $5.3$], since finite irretractable cycle sets with semiregular displacement group are latin.
@@
If $X$ is an irretractable cycle set, by a standard calculation we have that if either 1)1)1 ) or 2)2)2 ) hold, then $X$ is latin. Hence, the statement follows by [9, Theorem 5.3]. ∎
@@
If $(X,\cdotp)$ is an irretractable cycle set with regular displacement group, by 33 we can represent it as a cycle set right linear over a suitable group $(G,\circ)$ and suitable maps $\phi$ and $\psi$ as in 5. The theory contained in [9, Section 5] gives a method to find the operation $\circ$ and the maps $\phi$ and $\psi$, even if these maps are not uniquely determined. Indeed, if we choose two elements $e,f\in X$, then $X$ has a group structure given by $x\circ_{e,f}y:=\delta_{e}^{-1}(x)\cdotp\sigma_{f}^{-1}(y)$ for all $x,y\in X$ , where $f\cdotp e$ is the neutral element. In this way, $\phi_{e,f}(x):=\delta_{e}(x)\circ_{e,f}\sigma_{f}(e\cdotp f)$ and $\psi_{e,f}(x):=\sigma_{f}(e\cdotp f)^{-1}\circ_{e,f}\sigma_{f}(x)$ for all $x\in X$. The displacement group $\mathrm{Dis}(X)$ is isomorphic to $(X,\circ_{e,f})$, and an explicit isomorphism is given by $\rho_{e,f}:(X,\circ_{e,f})\rightarrow\mathrm{Dis}(X)$, $x\mapsto\sigma_{\delta_{e}^{-1}(x)}\sigma_{f}^{-1}$ for all $x\in X$. In this way, the map $\psi_{e,f}$ induces an automorphism of $\mathrm{Dis}(X)$ given by $\overline{\psi_{e,f}}(\sigma_{\delta_{e}^{-1}(x)}\sigma_{f}^{-1}):=\sigma_{\delta_{e}^{-1}(\psi_{e,f}(x))}\sigma_{f}^{-1}$ and the map $\phi_{e,f}$ induces a bijection given by $\overline{\phi_{e,f}}(\sigma_{\delta_{e}^{-1}(x)}\sigma_{f}^{-1}):=\sigma_{\delta_{e}^{-1}(\phi_{e,f}(x))}\sigma_{f}^{-1}$, for all $x\in X$. Therefore, under the correspondence $\rho_{e,f}$, the underlying group $(X,\circ_{e,f})$ can be identified with the displacement group and without loss of generality we can say that $X$ is right linear over $\mathrm{Dis}(X)$.
@@
By 33, $X$ can be identified with a cycle set right linear over a group $(G,\circ)$ and given by $x\cdotp y:=\phi(x)\circ\psi(y)$ for all $x,y\in G$, for a suitable maps $\phi$ and $\psi$. By 31 the congruences of $X$ are in one-to-one correspondence with normal subgroups $H$ of $G$ that are invariant under $\phi$ and $\psi$. Since $\mathrm{Dis}(X)$ acts on $X$ by left translations (respect to the operation $\circ$), the equivalence classes of the congruence induced by a subgroup $H$ of $G$ correspond to the orbits respect to the action of the subgroup $T_{H}$ of $\mathrm{Dis}(X)$ given by the translations by the elements of $H$. Therefore, since $X$ is irretractable, by [3, Lemma 2.1] and [35, Theorem 1] $T_{H}$ is an ideal of $\mathcal{G}(X)$. Therefore, every congruence comes from a suitable ideal of $\mathcal{G}(X)$ contained in $\mathrm{Dis}(X)$. Since $\mathrm{Dis}(X)$ acts by left translation with respect to the operation $\circ$, different ideals contained in $\mathrm{Dis}(X)$ give rise to different congruences, hence the statement follows. ∎
@@ 5.1 Simple left quasigroups
From now on, a cycle set will be called simple if its underlying left quasigroup is simple. As it is well-known that if $X$ is an indecomposable cycle set then the ideal $\mathcal{G}(X)^{2}$ of $\mathcal{G}(X)$ coincides with $\mathrm{Dis}(X)$ (see [11, Lemma 11]), we can use our results to recover a characterization of simple cycle sets. Since the unique simple decomposable cycle set has size $2$, we decide to add the condition $|X|>2$ and isolate this easy case.
@@
If $X$ is simple and is not the trivial cycle set of prime size, by [17, Proposition 4.1] it is irretractable, hence this implication follows by 20, 38 and the equality $\mathcal{G}(X)^{2}=\mathrm{Dis}(X)$. Conversely, if $X$ is the trivial indecomposable of prime size, then it is simple by [12, Lemma 1]. Now, suppose that $X$ is irretractable and $\mathcal{G}(X)^{2}$ is a minimal ideal of $\mathcal{G}(X)$ that acts transitively on $X$. Then $X$ is indecomposable and hence we have $\mathcal{G}(X)^{2}=\mathrm{Dis}(X)$. Therefore, the thesis follows by 20 and 38. ∎
@@ 5.2 Simple cycle sets with nilpotent displacement group
If $H$ is a non-trivial normal Hall subgroup of $\mathrm{Dis}(X)$, then by [4, Proposition 4.2.1] it is an ideal of $\mathrm{Dis}(X)$. Moreover, $H$ is a characteristic subgroup of $(\mathrm{Dis}(X),+)$ and $(\mathrm{Dis}(X),\circ)$, therefore it is an ideal of $\mathcal{G}(X)$. ∎
@@
1) $\Rightarrow$ 3) Suppose that $X$ is a simple cycle set with cyclic displacement group and such that $|X|$ is not a prime number. Then, $X$ must be irretractable and hence $\mathrm{Dis}(X)$ acts transitively on $X$, therefore by 41 we have $|X|=|\mathrm{Dis}(X)|=p^{k}$ for some prime number $p$. Moreover, by an inspection of the small cycle sets, $|X|\neq 4$. Now, let $m$ be a divisor of $|X|$ such that $|X|/m=p$ for a prime number $p$. By the results containted in [36] and [2, Proposition 5.4], $\mathrm{Dis}(X)$ is a left brace with cyclic additive and multiplicative group. Moreover, the set $H:=p\mathrm{Dis}(X)$ is an ideal of $\mathrm{Dis}(X)$ of size $m$, and is a characteristic subgroup of $(\mathrm{Dis}(X),+)$ and $(\mathrm{Dis}(X),\circ)$, hence it is an ideal of $\mathcal{G}(X)$. Therefore, $p\mathrm{Dis}(X)$ induces a non-trivial congruence on $X$, a contradiction. Then, $|X|$ must be a prime number, and by [17, Proposition 4.1] $|X|=2$ or $X$ is indecomposable. If the first case, clearly $\mathrm{Dis}(X)=\{id_{X}\}$ by a standard calculation. In the second case the conclusion is the same by [25, Theorem 2.13].
3) $\Rightarrow$ 2) If $X$ is a simple cycle set of prime size, we have that $\mathrm{Dis}(X)$ is the trivial group by [25, Theorem 2.13]